) Solve the system by substitution.
9x + 6y= -21
8x – 6y= -30
What is the solution of the system? Select the correct choice below and, if necessary, fill in the answer box within your choice.
a. The solution is ? (simplify your answer. Type an ordered pair. If there are infinitely many solutions, type an expression involving x for each coordinate of the ordered pair where x represents all real numbers.)
b. There is no solution.

Answer by mananth(16949)

9 x + 6 y = -21 .............1
-6 y = -30 -8 x ---------2
/ -6 =
y = 5 + 1.33 x
Substitute y in (1)
9 x + 6 ( 5 + 1.33 x ) = -21
9 x + 30 + 8 x = -21
9 x + 8 x = -21 -30
17 x = -51
/ 17
x= -3
Plug the value of x in (1)
9 x + 6 y = -21
9 * -3 + 6 y = -21
-27 + 6 y = -21
6 y = 6
/ 6
y= 1
(-3,1) is the ordered pair